OpenStudy (anonymous):

x 3 ----- = ------ x+2 x-2

OpenStudy (anonymous):

x/x+2=3/x-2

OpenStudy (anonymous):

^ what?

OpenStudy (anonymous):

That's the problem you posed rewritten in a conventional manner.

OpenStudy (anonymous):

\[\frac{ x }{ x+2 } = \frac{ 3 }{ x-2 }\]

OpenStudy (anonymous):

Yup. Same thing.